OpenStudy (anonymous):

-1/3x + 2/x = 5/(x+2) Find the value of x.

OpenStudy (pfenn1):

\[-\left( 1 \over 3x \right)+\left( 2 \over x \right)= \left( 5 \over x+2 \right)\] What would be the LCD?

OpenStudy (anonymous):

That's what I'm stuck on.

OpenStudy (anonymous):

Please help.

OpenStudy (pfenn1):

Okay. Let's just try 3x. What the equation be if you multiplied both sides by 3x?

OpenStudy (pfenn1):

\[-\left( 3x \over 3x \right)+\left( 6x \over x \right)= \left( 15x \over x+2 \right)\]\[-1+6=\left( 15x \over x+2 \right)\]\[5(x+2)=15x\]\[x+2=3x\]

OpenStudy (anonymous):

So x = 1?

OpenStudy (anonymous):

Oh, I plugged it in, it works.

OpenStudy (pfenn1):

Yes!

OpenStudy (anonymous):

Thank you! :D